OpenStudy (anonymous):

What's the square root of 72

OpenStudy (anonymous):

8.5

OpenStudy (anonymous):

or 8.48528137424

OpenStudy (anonymous):

Go on wolfram alpha and it automatically answers your questions for math

OpenStudy (anonymous):

Expressed as a surd in its simplest form, sqrt(7.2) = 6

OpenStudy (anonymous):

First, find out what numbers can multiply to equal 72 √72 = √9*8 Then find other numbers which can be squarerooted and will get you a whole number = √9*4*2 = √9 * √4 * √2 = 3* 2 * √2 =6√2